An election is accelerated from rest through a potential difference of 15V. What is the change in electric potential energy of the electron?

1
Expert's answer
2022-02-24T09:05:10-0500

The change in electric potential energy of the electron

"\Delta E_p=e\Delta V=1.6*10^{-19}*15=2.4*10^{-18}\:\rm J"